A boat that can move 3.0 m/s in still water is on a river that is flowing 1.0m/s (S) . What is the velocity of the boat relative to the shore when the boat is heading south?
1
Expert's answer
2013-03-13T08:09:50-0400
Velocity-addition formula:
if a boat is moving relative to the water velocity v and water is on a river that is flowing with velocity u,
then the velocity of the boat relative to the shore equals the vector sum:
s = v + u
In our case river is flowing to the south and boat is heading south, therefore:
s = v + u
s = 1m/s + 3 m/s = 4 m/s
Answer: the velocity of the boat relative to the shore equals 4 m/s